# Poro 34B and the Blessing of Multilinguality

### v2

@@ 4.2 Perplexity
Table 2: Average benchmark results for Finnish, English and code for Poro 34B and selected reference models
@@ 4.3 Benchmark results
The overall results of the benchmark evaluations are summarized in Table 2 and detailed in Appendix A.2. We find that Poro 34B is the best-performing model for Finnish in this comparison, substantially outperforming the best previously introduced monolingual Finnish model. We further analyzed the progression of the Finnish capabilities by evaluating Poro 34B checkpoints at 10% intervals on FIN-bench. These results are summarized in Figure 3. Interestingly, the model outperforms the best FinGPT model already after 100B tokens of training (10%) despite the relatively small proportion of Finnish in the Poro 34B data and the fact that the FinGPT models were trained on 300B tokens in total. These results indicate that our limited multilingual approach is effective for creating stronger models for Finnish than possible through monolingual training and demonstrate that the model is benefiting substantially from its training data in other languages when tested on Finnish tasks.
For English, we find that the model achieves broadly comparable results to the MPT 30B and Falcon 40B models, both of which were trained for 1T tokens of predominantly English data. This indicates that the limited multilingual training approach has not notably detracted from the English capabilities of the model. The best-performing open model in this comparison is Llama 33B, which was trained for longer (1.4T tokens), also predominantly on English data. We find that Poro 34B is nevertheless a capable model in its class also for English, despite not optimizing specifically for English performance. The programming language benchmarks indicate that Poro 34B is more capable on code than the other natural language-focused models, while the code-focused StarCoder model clearly outperforms all of the other models. We attribute the relatively high performance of Poro 34B on code to the comparatively large proportion of the training data dedicated to code. As with English, we consider the performance of the model on code a positive addition to its capabilities, even though code generation was not a primary goal in creating the model.
Finally, we note a surprising finding arising from the Finnish evaluation: two of the larger English-focused models (Llama 33B and MPT 30B) score higher than the previously introduced smaller monolingual Finnish models on the FIN-bench benchmark. While FIN-bench tasks are in Finnish, the benchmark consists of multiple-choice rather than generation tasks, has been produced in substantial part through translation from English, and includes tasks with little emphasis on natural language (e.g. arithmetic). We hypothesize that the comparatively high performance of the English-focused models on this benchmark might not indicate that they can generate fluent Finnish, which also calls the Finnish proficiency of Poro 34B into question. We study this question specifically in the following section.
@@ 4.4 Open-ended generation
To compare the ability of Poro 34B to generate coherent and grammatically correct texts in Finnish compared to other models, we prompted the models with short Finnish texts and evaluated the output using an LLM-as-a-judge approach (Zheng et al., 2023). To assure the prompt texts were not part of the training data of any of the models, we collected 50 recent news headlines from the Finnish broadcaster Yle [^11] https://yle.fi/. Headlines were collected on 11-12 January 2024. and 50 theses titles published in 2024 from an online archive [^12] https://www.finna.fi/. We prompted the models with the phrase Seuraava teksti on kirjoitettu suomeksi: (Eng.  The following text is written in Finnish:) followed by the title. We repeated the generation for each title five times, giving us 250 news and 250 academic output texts to evaluate. For every text from Poro 34B and the corresponding text from a baseline model, we ask GPT-4, as our judge model, to pick a preference based on grammatical correctness, coherence, and usage of vocabulary using a prompt adapted from that used in MT-bench (Zheng et al., 2023) (the prompt is provided in Appendix A.3). We query the judge twice for each pair of texts, switching the positions of the texts to eliminate any potential impact of ordering preference. We consider the pair a win for a model if the judge picks the text from that model twice, and a tie if the judge changes its preference when the order is switched.
FinGPT 8B FinGPT 13B Llama 33B MPT 30B Falcon 40B
News 52.0 / 12.4 53.2 / 13.2 2.0 / 8.4 10.8 / 11.6 0.8 / 6.4
Academic 46.0 / 10.0 50.0 / 10.4 0.8 / 10.4 21.2 / 7.2 0.4 / 10.0
Table 3: Win / tie rates (%) of reference models against Poro 34B in open-ended Finnish text generation.
The results of the experiments are summarized in Table 3. We find that the multilingual Poro 34B is roughly on par with the monolingual FinGPT models, performing competitively on academic text generation while falling narrowly behind on news text generation. By contrast, the English-focused models are clearly worse in generating Finnish text, with the best-performing model (MPT 30B on academic text) being preferred in only 21% of cases. The most striking contrast between the benchmark and open generation results is for Llama 33B, which scored higher than FinGPT on FIN-bench, but is preferred only in 2% of cases to Poro 34B in generation, where Poro 34B and FinGPT show roughly equal capabilities. These results support our hypothesis that multiple-choice benchmarks created (in substantial part) through translation of English benchmarks may be a poor measure of the specific language capabilities of large models. Unlike the English language models, Poro 34B continues to show strong performance here, showing that it is indeed both capable in general, and fluent in Finnish.
@@ 4.5 Translation
Table 4: spBLEU on the Flores-101 devtest and Tatoeba challenge test sets. Flores-101 results except for OPUS-MT and Poro 34B are from Zhu et al. (2023).
General-purpose language models have shown promising results on translation benchmarks on multiple languages (Vilar et al., 2023; Garcia et al., 2023; Alves et al., 2024). Following Zhu et al. (2023), we evaluated Poro 34B for English to Finnish translation and vice-versa on the first 100 sentences of the Flores-101 test data by prompting the model with eight translation examples sampled randomly from the development set, formatting the examples simply as <src>=<trg>. We further evaluated Poro 34B and three strong open-source translation models on the Tatoeba Challenge test set with more than 11,000 sentences: OPUS-MT (Tiedemann & Thottingal, 2020), NLLB-1.3B (Costa-jussà et al., 2022), and M2M-100-12B (Fan et al., 2021) [^13] We did not evaluate closed models such as ChatGPT on Tatoeba because of the cost of API access.. We used the standard SentencePiece BLEU (spBLEU) as our metric. The results of both evaluations are shown in Table 4.141414We attempted to reproduce some of the Flores-101 results reported by Zhu et al. (2023) and obtained a slightly higher result for GPT-4 in Eng-Fin translation (37.5 instead of 35.33) and slightly lower results for M2M-12B and NLLB-1.3B (31.4 and 26.6, respectively). For the sake of consistency, we present the results from that study without modification. These results demonstrate that Poro 34B is a remarkably strong translator, outperforming not only dedicated open-source translation models but even Google Translate, and scoring roughly on par with GPT-4 in this evaluation. We attribute this result to the combination of strong Finnish and English capabilities and the inclusion of a comparatively large number of translation examples in the pretraining data.
@@ 5 Discussion and conclusions
In this study, we have considered the challenges that the availability of data poses for pretraining large generative models for smaller languages and explored a limited multilingual approach to create Poro 34B, a 34B-parameter model trained on 1T tokens of Finnish, English, and code, including 8B tokens of Finnish-English translation pairs. We thoroughly evaluated the model and found it to substantially advance over the performance of existing models for Finnish while also performing competitively in its class of open models for English and code generation, as well as achieving remarkably good results in translation tasks. We additionally observed some limitations of benchmarks with tasks translated from English in measuring the specific capability of models in Finnish, which diverged substantially from model-as-judge evaluation of generation fluency.
Our model architecture and the Finnish datasets included follow those of the FinGPT family of monolingual Finnish models, which were limited by the available Finnish training data. The superior performance of our model in Finnish evaluations demonstrates that multilingual training can lift such limitations, allowing further scaling of models focused on smaller languages. In future work, we hope to explore this effect more systematically to answer some of the many questions that remain open regarding the training of large generative models for smaller languages, including the impacts of covering multiple smaller languages and the effect of the size of data available in the target languages.
@@
Reproducibility
It has been our aim throughout this work to release Poro 34B fully openly, including model weights, pretraining configuration, the pretraining and evaluation data, and all associated scripts and tools. We provide here additional details of these to facilitate accurate reproduction of our work.
The identifiers of the pretraining data sources are provided below in Table 5.
@@
Sequence length 2048 Min. learning rate 1e-5
@@
Position embedding ALiBi Weight decay 1e-1
@@
The model weights and a model card providing further details regarding the model and its pretraining are available from https://huggingface.co/LumiOpen/Poro-34B.
The model pretraining was performed using the Megatron-DeepSpeed fork https://github.com/TurkuNLP/Megatron-DeepSpeed.
All scripts and tools introduced specifically for this work are made available under open licenses at https://huggingface.co/LumiOpen/Poro-34B.
Ethics statement
Acknowledgements
Appendix A Appendix
A.1 Hardware
Poro 34B was trained on the LUMI-G GPU partition of the LUMI supercomputer, located in Finland. LUMI is at the time of this writing the fastest supercomputer in Europe, and the 5th fastest in the world (https://www.top500.org/). LUMI is also ranked 7th greenest by the Green500 list (https://www.top500.org/lists/green500/).
@@ A.2 Detailed benchmark results
Tables 7,  8 and 9 show the detailed benchmark results for Finnish, English, and code.
Table 7: FIN-Bench Finnish benchmark results
@@
Table 8: English benchmark results
@@
Table 9: Code benchmark results
@@
A.3 Judge prompt to GPT-4 for judging Finnish text
Please act as an impartial judge and evaluate the quality of two Finnish texts. Your evaluation should consider these three factors: 1. Vocabulary Usage: Wide range of vocabulary used effectively, including more specialized terms. 2. Grammatical Correctness: Strong grammatical skills; errors are rare and minor. 3. Coherence: The text is well-structured and coherent throughout.
Begin your evaluation by comparing the two texts and provide a short explanation. Avoid any position biases and ensure that the order in which the texts were presented does not influence your decision. Be as objective as possible. You are not allowed to declare a tie. After providing your explanation, output your final verdict by strictly following this format: "[[A]]" if Text A is better, "[[B]]" if Text B is better.

### v3

@@ 4.2 Perplexity
Table 2: Average benchmark results for Finnish, English and code for Poro 34B and selected reference models.
@@ 4.3 Benchmark results
The overall results of the benchmark evaluations are summarized in Table 2 and detailed in Appendix A.2. We find that Poro 34B is the best-performing model for Finnish in this comparison, substantially outperforming the best previously introduced monolingual Finnish model. We further analyzed the progression of the Finnish capabilities by evaluating Poro 34B checkpoints at 10% intervals on FIN-bench. These results are summarized in Figure 3. Interestingly, the model outperforms the best FinGPT model already after 100B tokens of training (10%) despite the relatively small proportion of Finnish in the Poro 34B data and the fact that the FinGPT models were trained on 300B tokens in total. These results indicate that our limited multilingual approach is effective for creating stronger models for Finnish than possible through monolingual training and demonstrate that the model is benefiting substantially from its training data in other languages even when tested on Finnish tasks.
For English, we find that the model achieves broadly comparable results to the MPT 30B and Falcon 40B models, both of which were trained for 1T tokens of predominantly English data. This indicates that the limited multilingual training approach has not notably detracted from the English capabilities of the model. The best-performing open model in this comparison is Llama 33B, which was trained for longer (1.4T tokens), also predominantly on English data. We find that Poro 34B is nevertheless a capable model in its class also for English, despite not optimizing specifically for English performance. The programming language benchmarks indicate that Poro 34B is more capable on code than the other natural language-focused models, while the code-focused StarCoder model clearly outperforms all of the other models. We attribute the relatively high performance of Poro 34B on code to the comparatively large proportion of the training data dedicated to code. As with English, we consider the performance of the model on code a positive addition even though code generation was not a primary goal in creating the model.
Figure 4: Win counts of reference models against Poro 34B on Finnish MT-Bench as judged by GPT-4 Turbo.
Finally, we note a surprising finding arising from the Finnish evaluation: two of the larger English-focused models (Llama 33B and MPT 30B) score higher than the previously introduced smaller monolingual Finnish models on the FIN-bench benchmark. While FIN-bench tasks are in Finnish, the benchmark consists of multiple-choice rather than generation tasks, has been produced in substantial part through translation from English, and includes tasks with little emphasis on natural language (esp. arithmetic). We hypothesize that the comparatively high performance of the English-focused models on this benchmark might not indicate that they can generate fluent Finnish, which also calls the Finnish proficiency of Poro 34B into question. We study this question specifically in the following section.
@@ 4.4 Open-ended generation
To assess the ability of the models to generate coherent and grammatically correct Finnish, we create a Finnish version of MT-Bench Zheng et al. (2023), a benchmark for open-ended conversations that uses LLM-as-a-judge evaluation. We excluded math and coding questions to focus specifically on the natural language generation capabilities of the models. To create the benchmark, we initially translated the questions into Finnish using DeepL, [^11] https://www.deepl.com and the translations were then manually corrected by native Finnish speakers to create the final evaluation dataset. To evaluate base models using the data, we similarly translated and corrected the few-shot URIAL prompt Lin et al. (2024). 121212 We did not modify the judge prompts as previous work has found that keeping the prompt in English produces better results Ahuja et al. (2023). We use pairwise judging to compare between Poro 34B and the competing models’ responses and use GPT-4 Turbo as the judge model.
To assess the reliability of the model as a judge and provide further insight into the quality of the generations, we additionally set up an annotation platform where two native Finnish speakers were asked to pick a preference between a response generated by Poro 34B and a competing model. [^13] We did not separately compensate the human judges as they are co-authors of this paper. The judges are given the same judging prompt as GPT. The model names are hidden from the judges, and we randomly select the position of each response in every response pair to account for positional bias.
We found that the two human judges highly agree with each other, picking the same winner 88.8% of the time, and found an even higher agreement between GPT and each human judge: 91.6% between annotator 1 and GPT and 89.5% between annotator 2 and GPT. Figure 4 shows the win counts of the reference models against Poro 34B as judged by GPT-4 Turbo.
In manual analysis after the initial annotation, we found that the FinGPT models often struggled with the few-shot format, failing to follow questions or only giving short, minimal answers, while Poro 34B was better able to comply with questions and given requirements, such as listing a specified number of items. However, we found that Poro 34B also often hallucinated and did not follow all instructions, and we would not consider its responses to be at a level of consistency and quality required for user-facing applications, which is not an unexpected result given that it is a base model not specifically fine-tuned or otherwise aligned for such use. Despite outperforming FinGPT models on the FIN-bench benchmark, The English-focused models appeared to be unfit for Finnish generation: their generations had the surface appearance of Finnish text but were largely nonsensical and incoherent. This result underlines the need to include multiple perspectives when evaluating models: a high score on a multiple-choice benchmark may not indicate practical capability to generate coherent text in a language.
We make the Finnish MT-Bench available under an open license and provide the model generations at https://github.com/LumiOpen/FastChat/tree/main/fastchat/llm_judge.
@@ 4.5 Translation
General-purpose language models have shown promising results on translation benchmarks on multiple languages (Vilar et al., 2023; Garcia et al., 2023; Alves et al., 2024). Following Zhu et al. (2023), we evaluated Poro 34B for English to Finnish translation and vice versa on the first 100 sentences of the Flores-101 test data by prompting the model with eight translation examples sampled randomly from the development set, formatting the examples simply as <src>=<trg>. We further evaluated Poro 34B and three strong open-source translation models on the Tatoeba test set with more than 11,000 sentences: OPUS-MT (Tiedemann and Thottingal, 2020), NLLB-1.3B (Costa-jussà et al., 2022), and M2M-100-12B (Fan et al., 2021) [^14] We did not evaluate the GPT models and Google Translate on Tatoeba because of the associated API costs.. We used the standard SentencePiece BLEU (spBLEU) as our metric. The results of both evaluations are shown in Table 3.151515We attempted to reproduce some of the Flores-101 results reported by Zhu et al. (2023) and obtained a slightly higher result for GPT-4 in Eng-Fin translation (37.5 instead of 35.33) and slightly lower results for M2M-12B and NLLB-1.3B (31.4 and 26.6, respectively). For the sake of consistency, we present the results from that study without modification. These results demonstrate that Poro 34B is a remarkably strong translator, outperforming not only dedicated open-source translation models but even Google Translate, and scoring roughly on par with GPT-4 in this evaluation. We attribute this result to the combination of strong Finnish and English capabilities and the inclusion of a comparatively large number of translation examples in the pretraining data.
@@
Table 3: spBLEU on the Flores-101 devtest and Tatoeba test sets. Flores-101 results except for OPUS-MT and Poro 34B are from Zhu et al. (2023).
@@ 5 Discussion and conclusions
In this study, we have considered the challenges that the availability of data poses for pretraining large generative models for smaller languages and explored a limited multilingual approach to create Poro 34B, a 34B-parameter model trained on 1T tokens of Finnish, English, and code, including 8B tokens of Finnish-English translation pairs. We thoroughly evaluated the model and found it to substantially advance over the performance of existing models for Finnish while also performing competitively in its class of open models for English and code generation, as well as achieving remarkably good results in translation tasks. Two human judges and GPT-4 Turbo found the texts generated by Poro 34B to be superior to the competing models.
Our model architecture and the Finnish datasets included follow those of the FinGPT family of monolingual Finnish models, which were constrained by the available Finnish training data. The superior performance of our model in Finnish evaluations demonstrates that multilingual training can lift such limitations, allowing further scaling of models focused on smaller languages. In future work, we hope to explore this effect more systematically to answer some of the many questions that remain open regarding the training of large generative models for smaller languages, including the impacts of covering multiple smaller languages and the effect of the size of data available in the target languages.
@@
Limitations
Our study applies a pretraining recipe that combines insights on effective multilingual and data-constrained model training from a variety of previous studies. While the findings of these studies are supported by a broad range of relevant experimental results, we did not have the resources to perform separate ablation experiments specifically assessing the impact that various parts of our combined pretraining recipe (e.g., four repetitions of target language data and the inclusion of a translation signal) have on the resulting model. Thus, while we believe that our results demonstrate the pretraining recipe to be effective for creating state-of-the-art models for data-constrained languages, our work is limited in leaving many questions open regarding specific choices that form part of that recipe.
Poro 34B is a base model and as such has not been aligned to follow instructions and engage in conversations. It has not been evaluated on safety and toxicity benchmarks. As we have noted in our language generation evaluation, Poro 34B does not adequately follow instructions and has the tendency to generate texts with hallucinations. Further research is needed to improve the model in terms of factuality, safety, and alignment in English and Finnish. We encourage developers using Poro 34B to be aware of the potential risks associated with LLMs such as non-factual outputs, harmful language, and perpetuation of biases and stereotypes. We recommend that developers finetune Poro 34B to meet their specific needs and codes of conduct.
Ethical considerations
Acknowledgements
Appendix A Appendix
A.1 Training details
It has been our aim throughout this work to release Poro 34B fully openly, including model weights, pretraining configuration, the pretraining and evaluation data, and all associated scripts and tools. We provide here additional details of these to facilitate accurate reproduction of our work. The pretraining data sources are detailed in Table 4, and the model and pretraining hyperparameters in Table 5.
@@ A.1 Training details
Sequence length 2048 Min. learning rate 2e-5
@@
Position embedding ALiBi Weight decay 2e-5
@@ A.2 Detailed benchmark results
Tables 67, and 8 show the detailed benchmark results for Finnish, English, and code.
@@
Table 6: FIN-Bench Finnish benchmark results
@@
Table 7: English benchmark results
@@
Table 8: Code benchmark results
A.3 Hardware
Poro 34B was trained on the LUMI-G GPU partition of the LUMI supercomputer, located in Finland. LUMI is, at the time of this writing, the third fastest supercomputer in Europe, and the 8th fastest in the world (https://www.top500.org/). LUMI is also ranked 7th greenest by the Green500 list (https://www.top500.org/lists/green500/).